OpenStudy (anonymous):

The area of a circle with radius r is given by the formula A=pie r^2. If a circle has area 36pie(sign) square inches, what is the radius?

OpenStudy (saifoo.khan):

\[Area = \pi * r^2\]\[36 \pi = \pi *r^2\]